# Contextual Embedding Learning to Enhance 2D Networks for Volumetric Image Segmentation

### v1

@@ Abstract
The segmentation of organs in volumetric medical images plays an important role in computer-aided diagnosis and treatment/surgery planning. Conventional 2D convolutional neural networks (CNNs) can hardly exploit the spatial correlation of volumetric data. Current 3D CNNs have the advantage to extract more powerful volumetric representations but they usually suffer from occupying excessive memory and computation nevertheless. In this study we aim to enhance the 2D networks with contextual information for better volumetric image segmentation. Accordingly, we propose a contextual embedding learning approach to facilitate 2D CNNs capturing spatial information properly. Our approach leverages the learned embedding and the slice-wisely neighboring matching as a soft cue to guide the network. In such a way, the contextual information can be transferred slice-by-slice thus boosting the volumetric representation of the network. Experiments on challenging prostate MRI dataset (PROMISE12) and abdominal CT dataset (CHAOS) show that our contextual embedding learning can effectively leverage the inter-slice context and improve segmentation performance. The proposed approach is a plug-and-play, and memory-efficient solution to enhance the 2D networks for volumetric segmentation. The code will be publicly available.
@@
journal: arXiv
@@ 1 Introduction
The segmentation of organs in volumetric medical images (e.g., computed tomography (CT), magnetic resonance imaging (MRI), 3D ultrasound) enables quantitative analysis of anatomical parameters such as shape, boundary, volume, etc [31, 22]. Also, the segmentation is often the very first step in the workflow of other medical image computing tasks, such as computer-assisted diagnosis [11, 40], and surface-based registration [10, 41]. However, it is tedious and time-consuming for the clinicians to manually delineate the organ’s contour. Therefore, the automatic segmentation approaches are highly needed to satisfy the constantly increasing clinical demand.
@@
With the promising capability of deep learning techniques in the society of computer vision and pattern recognition [18], convolutional neural networks (CNNs) have also become a primary option for medical image segmentation. Various CNN architectures have achieved excellent segmentation results in many medical applications. One of the most well-known CNN architectures for medical image segmentation is the U-net [30], which adopts the skip connections to effectively aggregate low-level and high-level features. The U-net has now been the benchmark for lots of segmentation tasks and based on which many 2D variants have been devised to target specific applications. Chen et al. [3] proposed a multi-task learning network for gland segmentation from histopathological images, which explores multi-level contextual features from the auxiliary classifier on top of the U-net structure. Ibtehaz et al. [15] took inspiration from Inception blocks [33] and proposed MultiRes blocks to enhance U-Net with the ability of multi-resolutional analysis. Zhou et al. [47] re-designed the skip connections and proposed a UNet++ which is a deeply-supervised encoder-decoder architecture with nested and dense skip connections. Jha et al. [16] proposed ResUNet++, which took advantage of residual blocks, squeeze and excitation blocks, and strous spatial pyramidal pooling (ASPP) to segment the colorectal polyps. However, directly employing 2D networks to deal with 3D images in a slice-by-slice way may result in discontinuous or inaccurate segmentation in 3D space (see the conventional 2D results in Fig. 1), due to the limited volumetric representation capability of 2D learning [19]. To better extract spatial information, some 2.5D approaches have been proposed [29, 36]. In general, these approaches still employ 2D convolutional kernels, whereas extracting features from three orthogonal views, i.e., transversal, coronal, and sagittal planes. Although 2.5D approaches could slightly improve the 2D networks’ results, they still may not well exploit the original volumetric images [17]. Moreover, the three branches of 2.5D networks would cause a rapid increase in parameters and memory usage.
@@
To better empower the networks with richer volumetric information, Çiçek et al. [7] proposed 3D U-net to extend U-net architecture directly processing 3D images. However, the high memory consumption of 3D U-net would limit the depth of networks. Milletari et al. [24] proposed another 3D derivation of U-net architecture, named V-net, applying residual connections to enable a deeper network. Due to the advantages of residual connections, several studies [2, 39, 42] further employed the residual design to boost the 3D networks. Furthermore, to better aggregate the multi-scale spatial features, various attention mechanisms have been added into the 3D CNNs [26, 37, 20, 14]. Although 3D CNNs are beneficial to extract more powerful volumetric features, they still suffer from occupying excessive memory and computation.
@@
It is worth noting that combining CNNs with Recurrent Neural Networks (RNNs), such as Long Short-Term Memory (LSTM) [12] and convolutional-LSTM [32], can obtain slice-wise contextual information. Chen et al. [4] combined a 2D U-net with bi-directional LSTM-RNNs to mine the intra-slice and inter-slice contexts for 3D image segmentation. Poudel et al. [28] devised a recurrent fully-convolutional network to learn inter-slice spatial dependences for cardiac MRI segmentation. Alom et al. [1] developed a recurrent residual U-net model (R2U-Net), which utilized the recurrent residual convolutional layers to ensure better feature representation for segmentation tasks. Novikov et al. [25] proposed Sensor3D to integrate bi-directional C-LSTMs into a U-net-like architecture to extract the inter-slice context sequentially, which achieved notable performance in CT segmentation. RNNs can extract the contextual relationship, yet they bring extra computational cost and memory usage.
@@ 2.4 Attention Merge Module
where $u$ denotes the prediction (i.e., the $P_{n^{\prime}}$ and $P_{n}$), $W$ and $H$ represent its dimension in width and height, respectively. In order to fully capture the dependency between predictions, we then employ a simple gating mechanism with a sigmoid activation ($\sigma$):
@@
Methods DSC (%) ASSD (mm) 95HD (mm)
V-Net [24] 85.12 $\pm$ 2.28* 1.95 $\pm$ 0.61* 6.13 $\pm$ 4.71
U-Net [30] 85.67 $\pm$ 6.89* 2.05 $\pm$ 2.36* 6.92 $\pm$ 10.88
3D U-Net [7] 86.03 $\pm$ 6.79 2.07 $\pm$ 2.08* 6.70 $\pm$ 11.17
U-Net+CE (Ours) 87.38 $\pm$ 4.42 1.55 $\pm$ 0.76 5.14 $\pm$ 4.01
DeepLabV3 [5] 83.37 $\pm$ 6.21* 2.13 $\pm$ 1.01* 7.34 $\pm$ 6.48*
3D DeepLabV3 84.86 $\pm$ 6.49 1.88 $\pm$ 0.82 4.83 $\pm$ 2.33
DeepLabV3+CE (Ours) 84.94 $\pm$ 5.80 1.94 $\pm$ 0.91 6.26 $\pm$ 5.14
@@ 3.1 Experimental Data and Pre-processing
The prostate MRI images from the PROMISE12 dataset were collected from four different centers, thus possessing diverse appearances (see Fig. 5). We used 50 MRI volumes to conduct five-fold cross validation. The intra- and inter-slice resolutions of these volumes range from 0.25 mm to 0.75 mm, and from 2.2 mm to 4.0 mm, respectively. All MRI volumes were resampled to a fixed resolution of 1$\times$1$\times$1.5 mm3. The N4 bias field correction [34] was conducted, and then the intensity values were normalized to zero mean and unit variance.
The liver CT volumes were from the challenging CHAOS dataset. Fig. 6 illustrates some example CT images, which show low tissue contrast and ambiguous organs’ boundaries. We used 40 CT volumes to conduct five-fold cross validation. All CT slices were resampled to the size of 256$\times$256. The intensity values were clipped to an interval of [-100, 200] HU, and then normalized to zero mean and unit variance. During training, contrast stretching, brightness stretching and elastic deformation were applied for data augmentation.
@@ 3.2 Implementation Details
The segmentation model was implemented on the open source platform Pytorch. The values of neighboring interval $l$, local patch size $k$, coefficient $r$ were empirically set to 1, 3, 64, respectively. Dice loss was utilized to train the model. Stochastic gradient descent (SGD) was used with batch size of 50, learning rate of 0.01, momentum of 0.9 and weight decay of 0.001 to control the learning progress of the network. The total number of training epochs was 300. All experiments were conducted on a single NVIDIA GeForce GTX 2080 GPU.
@@ 3.3 Evaluation Metrics
The metrics employed to quantitatively evaluate the segmentation accuracy included Dice similarity coefficient (DSC), average symmetric surface distance (ASSD), and 95% Hausdorff distance (95HD) [38]. All above metrics were calculated in 3D space. DSC measures the relative volumetric overlap between the predicted and ground-truth segmentations. ASSD determines the average distance between the surfaces of the predicted and ground-truth segmentations. 95HD is similar to ASSD but more sensitive to the localized disagreement as it determines the 95th percentile of all calculated Hausdorff distances. A better segmentation shall have smaller ASSD and 95HD, and larger value of DSC metric.
@@
3.4 Segmentation Accuracy
To demonstrate the efficacy of the proposed CE block, we applied it to enhance two popular 2D segmentation networks, i.e., U-Net [30] and DeepLabV3 [5], in a plug-and-play manner. We compared the CE-enhanced 2D networks with their 3D version, i.e., 3D U-Net [7] and 3D DeepLabV3. We also compared our method with two well-established models for prostate MRI segmentation (V-Net [24]) and liver CT segmentation (U-Net++ [47]), respectively.
@@ 3.4 Segmentation Accuracy
Table 1 reports the numerical results of all methods on the prostate MRI dataset. To investigate the statistical significance of the proposed method over compared methods on each metrics, the Wilcoxon signed-rank test was employed and also reported in Table 1. It can be observed that our CE-enhanced 2D networks can be regarded as significantly better than the conventional 2D networks on almost all evaluation metrics (except for the pair U-Net $vs$. U-Net$+$CE on 95HD). Furthermore, the $p$-values of CE-enhanced 2D networks $vs$. 3D networks on almost all metrics are beyond the 0.05 level, which indicates that the CE-enhanced 2D networks achieved similar performance as the 3D networks did. Fig. 7 visualizes the segmentation results in transversal, coronal and sagittal views, respectively. Apparently, our CE-enhanced networks obtained the most similar segmented boundaries to the ground truth. Compared with conventional 2D networks, our method generated more accurate and continuous results in volumetric view (see the coronal and sagittal views in Fig. 7). In general, the results shown in Table 1 and Fig. 7 prove the effectiveness of our contextual embedding strategy. Compared with conventional 2D networks, the proposed CE block leveraged the learned embedding and the slice-wisely neighboring matching as an internal guidance to transfer contextual information thus facilitating 2D networks to capture volumetric representation properly.
The quantitative results on the liver CT dataset are listed in Table 2. It can be observed that the CE-enhanced networks outperformed conventional 2D networks on all metrics. It is worth noting that the proposed CE block largely improved the 2D networks with respect to the metrics of ASSD and 95HD, corroborating that the CE block has strong capability to utilize the slice-wise context for the accurate volumetric segmentation. The statistical results in Table 2 further shows our method can make the segmentation performance of the 2D networks close to or even surpass that of the 3D networks. Figs. 8 visualizes the liver segmentation results in transversal, coronal and sagittal views, respectively. Our method can successfully infer the ambiguous boundaries and attain an accurate and smooth volumetric segmentation.
Methods DSC (%) ASSD (mm) 95HD (mm)
UNet++ [47] 94.23 $\pm$ 2.94* 0.95 $\pm$ 0.58* 4.89 $\pm$ 6.50*
U-Net [30] 94.13 $\pm$ 2.32* 1.84 $\pm$ 1.12* 8.62 $\pm$ 7.62*
3D U-Net [7] 95.67 $\pm$ 2.23* 2.31 $\pm$ 1.72* 2.31 $\pm$ 1.72
U-Net+CE (Ours) 96.49 $\pm$ 1.01 0.50 $\pm$ 0.17 1.77 $\pm$ 0.84
DeepLabV3 [5] 95.05 $\pm$ 4.85* 1.78 $\pm$ 2.24* 7.62 $\pm$ 14.83*
3D DeepLabV3 95.33 $\pm$ 1.44 1.49 $\pm$ 0.81* 5.83 $\pm$ 5.56*
DeepLabV3+CE (Ours) 95.59 $\pm$ 2.03 0.61 $\pm$ 0.37 1.95 $\pm$ 1.50
     Methods      ##\## Params      FLOPs
     UNet [30]      13.39 M      1092.31
     3D UNet [7]      40.15 M      1611.50
     U-Net+CE (Ours)      24.15 M      1466.12
     DeepLabV3 [5]      25.42 M      941.41
     3D DeepLabV3      74.72 M      1423.17
     DeepLabV3+CE (Ours)      37.40 M      1373.74

### v2

@@ Abstract
The segmentation of organs in volumetric medical images plays an important role in computer-aided diagnosis and treatment/surgery planning. Conventional 2D convolutional neural networks (CNNs) can hardly exploit the spatial correlation of volumetric data. Current 3D CNNs have the advantage to extract more powerful volumetric representations but they usually suffer from occupying excessive memory and computation nevertheless. In this study we aim to enhance the 2D networks with contextual information for better volumetric image segmentation. Accordingly, we propose a contextual embedding learning approach to facilitate 2D CNNs capturing spatial information properly. Our approach leverages the learned embedding and the slice-wisely neighboring matching as a soft cue to guide the network. In such a way, the contextual information can be transfered slice-by-slice thus boosting the volumetric representation of the network. Experiments on challenging prostate MRI dataset (PROMISE12) and abdominal CT dataset (CHAOS) show that our contextual embedding learning can effectively leverage the inter-slice context and improve segmentation performance. The proposed approach is a plug-and-play, and memory-efficient solution to enhance the 2D networks for volumetric segmentation. Our code is publicly available at https://github.com/JuliusWang-7/CE_Block.
@@
journal: Expert Systems With Applications
@@ 1 Introduction
The segmentation of organs in volumetric medical images (e.g., computed tomography (CT), magnetic resonance imaging (MRI), 3D ultrasound) enables quantitative analysis of anatomical parameters such as shape, boundary, volume, etc [38, 28, 53]. Also, the segmentation is often the very first step in the workflow of other medical image computing tasks, such as computer-assisted diagnosis [15, 49, 33], and surface-based registration [45, 12, 50]. However, it is tedious and time-consuming for the clinicians to manually delineate the organ’s contour. Therefore, the automatic segmentation approaches are highly needed to satisfy the constantly increasing clinical demand.
@@
With the promising capability of deep learning techniques in the society of computer vision and pattern recognition [24], convolutional neural networks (CNNs) have also become a primary option for medical image segmentation. Various CNN architectures have achieved excellent segmentation results in many medical applications. One of the most well-known CNN architectures for medical image segmentation is the U-net [37], which adopts the skip connections to effectively aggregate low-level and high-level features. The U-net has now been the benchmark for lots of segmentation tasks and based on which many 2D variants have been devised to target specific applications. Chen et al. [4] proposed a multi-task learning network for gland segmentation from histopathological images, which explores multi-level contextual features from the auxiliary classifier on top of the U-net structure. Ibtehaz et al. [20] took inspiration from Inception blocks [40] and proposed MultiRes blocks to enhance U-Net with the ability of multi-resolutional analysis. Zhou et al. [54] re-designed the skip connections and proposed a UNet++ which is a deeply-supervised encoder-decoder architecture with nested and dense skip connections. Jha et al. [21] proposed ResUNet++, which took advantage of residual blocks, squeeze and excitation blocks, and strous spatial pyramidal pooling (ASPP) to segment the colorectal polyps. Gao et al. [11] introduced UTNet, a hybrid Transformer architecture that blended self-attention mechanism with CNNs to enhance medical image segmentation. However, directly employing 2D networks to deal with 3D images in a slice-by-slice way may result in discontinuous or inaccurate segmentation in 3D space (see the conventional 2D results in Fig. 1), due to the limited volumetric representation capability of 2D learning [25]. To better extract spatial information, some 2.5D approaches have been proposed [36, 46, 43]. In general, these approaches still employ 2D convolutional kernels, whereas extracting features from three orthogonal views, i.e., transversal, coronal, and sagittal planes. Although 2.5D approaches could slightly improve the 2D networks’ results, they still may not well exploit the original volumetric images [22]. Moreover, the three branches of 2.5D networks would cause a rapid increase in parameters and memory usage.
@@
To better empower the networks with richer volumetric information, Çiçek et al. [8] proposed 3D U-net to extend U-net architecture directly processing 3D images. However, the high memory consumption of 3D U-net would limit the depth of networks. Milletari et al. [30] proposed another 3D derivation of U-net architecture, named V-net, applying residual connections to enable a deeper network. Due to the advantages of residual connections, several studies [3, 48, 51] further employed the residual design to boost the 3D networks. Furthermore, to better aggregate the multi-scale spatial features, various attention mechanisms have been added into the 3D CNNs [32, 47, 26, 18, 2, 52]. Hatamizadeh et al. [14] presented a novel architecture, called UNETR (UNET Transformers), which employed a Transformer as the encoder to acquire sequence representations of the input volume and captured global multi-scale information. Although 3D CNNs are beneficial to extract more powerful volumetric features, they still suffer from occupying excessive memory and computation.
@@
It is worth noting that combining CNNs with Recurrent Neural Networks (RNNs), such as Long Short-Term Memory (LSTM) [16] and convolutional-LSTM [39], can obtain slice-wise contextual information. Chen et al. [5] combined a 2D U-net with bi-directional LSTM-RNNs to mine the intra-slice and inter-slice contexts for 3D image segmentation. Poudel et al. [35] devised a recurrent fully-convolutional network to learn inter-slice spatial dependences for cardiac MRI segmentation. Alom et al. [1] developed a recurrent residual U-net model (R2U-Net), which utilized the recurrent residual convolutional layers to ensure better feature representation for segmentation tasks. Novikov et al. [31] proposed Sensor3D to integrate bi-directional C-LSTMs into a U-net-like architecture to extract the inter-slice context sequentially, which achieved notable performance in CT segmentation. Kang et al. [23] combined 3D CNN with ConvLSTM for fine segmentation, aiming to better capture contextual information between slices and thereby improving segmentation results. RNNs can extract the contextual relationship, yet they bring extra computational cost and memory usage.
@@ 2.4 Attention Merge Module
where $u$ denotes the concatenation of $P_{n}$ and $P_{n^{\prime}}$, $W$ and $H$ represent its dimension in width and height, respectively. In order to fully capture the dependency between predictions, we then employ a simple gating mechanism with a sigmoid activation ($\sigma$):
@@ 3.1 Experimental Data and Pre-processing
The prostate MRI images from the PROMISE12 dataset were collected from four different centers, thus possessing diverse appearances (see Fig. 5). We used 50 MRI volumes to conduct patient-wise five-fold cross validation. The intra- and inter-slice resolutions of these volumes ranged from 0.25 mm to 0.75 mm, and from 2.2 mm to 4.0 mm, respectively. Each 3D volume comprised between 15 and 54 slices. All MRI volumes were resampled to a fixed resolution of 1$\times$1$\times$1.5 mm3. The N4 bias field correction [42] was conducted, and then the intensity values were normalized to zero mean and unit variance.
The liver CT volumes were from the challenging CHAOS dataset. Fig. 6 illustrates some example CT images, which show low tissue contrast and ambiguous organs’ boundaries. We used 40 CT volumes to conduct patient-wise five-fold cross validation. All CT slices were resampled to the size of 256$\times$256. Each 3D volume comprised between 81 and 266 slices. The intensity values were clipped to an interval of [-100, 200] HU, and then normalized to zero mean and unit variance. During training, contrast stretching, brightness stretching and elastic deformation were applied for data augmentation.
Require: Image $S_{n}$, label $G_{n}$($n$ represents n-th slice of volumetric image)
Require: Backbone features $B_{n}$, backbone prediction $P_{n}$, slice-wisely neighboring matching prediction $P_{n^{\prime}}$, refined prediction $P_{n-CE}$, distance maps $D_{n}$
Require: Backbone model $f(\cdot)$ with parameters $\theta_{b}$
Require: Embedding space transform module $e(\cdot)$ with parameters $\theta_{e}$
Require: Slice-wisely neighboring matching module $s(\cdot)$ with parameters $\theta_{s}$
Require: Attention Merge Module $m(\cdot)$ with parameters $\theta_{m}$
Initialize: $\theta_{b}$, $\theta_{e}$, $\theta_{s}$ and $\theta_{m}$
for t in [1, num_epochs] do
       for i in [1, num_patients] do
[…]
            ($B_{0}$, …, $B_{n}$), ($P_{0}$, …, $P_{n}$) = $f(S_{0},...,S_{n})$,
[…]
            $E_{n}$ = $e(B_{n})$
[…]
            $D_{n}$ = $d(E_{n-l},E_{n},P_{n})$
[…]
            $P_{n^{\prime}}$ = $s(P_{n-l},D_{n},B_{n})$
[…]
            $P_{n-CE}$ = $m(P_{n},P_{n^{\prime}})$
            Compute Loss $\mathcal{L}_{DSC}$
            Update $\theta_{b}$, $\theta_{e}$, $\theta_{s}$ and $\theta_{m}$ using SGD
       end for
end for
return $\theta_{b}$, $\theta_{e}$, $\theta_{s}$ and $\theta_{m}$
Algorithm 1 Contextual Embedding Learning for Volumetric Image Segmentation using 2D Networks
Methods DSC (%) ASSD (mm) 95HD (mm) Sensitivity (%) Precision (%)
V-Net [30] 85.12 $\pm$ 2.28* 1.95 $\pm$ 0.61* 6.13 $\pm$ 4.71 91.22 $\pm$ 4.62 83.41 $\pm$ 8.27*
U-Net [37] 85.67 $\pm$ 6.89* 2.05 $\pm$ 2.36* 6.92 $\pm$ 10.88 85.94 $\pm$ 6.55* 85.05 $\pm$ 9.40*
3D U-Net [8] 86.03 $\pm$ 6.79 2.07 $\pm$ 2.08* 6.70 $\pm$ 11.17 89.26 $\pm$ 5.94 87.73 $\pm$ 7.04
U-Net+CE (Ours) 87.38 $\pm$ 4.42 1.55 $\pm$ 0.76 5.14 $\pm$ 4.01 88.50 $\pm$ 4.07 87.73 $\pm$ 8.58
DeepLabV3 [6] 83.37 $\pm$ 6.21* 2.13 $\pm$ 1.01* 7.34 $\pm$ 6.48* 80.52 $\pm$ 7.63 89.80 $\pm$ 9.19
3D DeepLabV3 84.86 $\pm$ 6.49 1.88 $\pm$ 0.82 4.83 $\pm$ 2.33 84.83 $\pm$ 6.93 84.09 $\pm$ 8.83
DeepLabV3+CE (Ours) 84.94 $\pm$ 5.80 1.94 $\pm$ 0.91 6.26 $\pm$ 5.14 82.31 $\pm$ 5.41 87.27 $\pm$ 9.35
Missformer [19] 82.81 $\pm$ 7.75* 2.13 $\pm$ 1.14 5.97 $\pm$ 3.25* 85.89 $\pm$ 6.55 82.22 $\pm$ 14.62
SwinUNETR [13] 87.74 $\pm$ 3.32* 1.31 $\pm$ 0.27* 4.11 $\pm$ 0.71 91.45 $\pm$ 3.22* 84.91 $\pm$ 7.60
Missformer+CE (Ours) 83.70 $\pm$ 6.39 2.19 $\pm$ 1.25 4.80 $\pm$ 3.79 86.00 $\pm$ 6.54 83.74 $\pm$ 13.47
@@ 3.2 Implementation Details
The segmentation model was implemented on the open source platform Pytorch. The values of neighboring interval $l$, local patch size $k$, coefficient $r$ were empirically set to 1, 3, 64, respectively. Dice loss was utilized to train the model:
$\mathcal{L}_{\mathcal{DSC}}=1-\frac{1}{C}\sum_{c=1}^{C}\frac{2\sum_{i=1}^{N}{y_{i,c}p_{i,c}}}{\sum_{i=1}^{N}{y_{i,c}+\sum_{i=1}^{N}p_{i,c}}},$ (6)
where $y_{i,c}$ employed a one-hot encoding scheme representing the ground truth labels, $p_{i,c}$ denoted the predicted values produced by the model for each class, and where indices i and c iterated over all pixels and classes, respectively. Stochastic gradient descent (SGD) was used with batch size of 50, learning rate of 0.01, momentum of 0.9 and weight decay of 0.001 to control the learning progress of the network. The total number of training epochs was 300. All experiments were conducted on a single NVIDIA GeForce GTX 2080 GPU.
The Algorithm 1 provides the pseudo-code of our proposed method.
@@ 3.3 Evaluation Metrics
The metrics employed to quantitatively evaluate the segmentation accuracy included Dice similarity coefficient (DSC), average symmetric surface distance (ASSD), 95% Hausdorff distance (95HD), sensitivity and precision [41]. All above metrics were calculated in 3D space. DSC is the primary metric used for assessing the segmentation accuracy, which measures the relative volumetric overlap between the predicted and ground-truth segmentations. ASSD determines the average distance between the surfaces of the predicted and ground-truth segmentations. 95HD is similar to ASSD but more sensitive to the localized disagreement as it determines the 95th percentile of all calculated Hausdorff distances. Sensitivity indicates the model’s capability to detect the foreground class in an image. Precision indicates the model’s capability to predict correct positive samples in an image. A better segmentation shall have smaller ASSD and 95HD, and larger values of DSC, precision and sensitivity.
3.4 Segmentation Accuracy
To demonstrate the efficacy of the proposed CE block, we applied it to enhance three popular 2D segmentation networks, i.e., U-Net [37], DeepLabV3 [6] and Missformer [19], in a plug-and-play manner. We compared the CE-enhanced 2D networks with their 3D version, i.e., 3D U-Net [8], 3D DeepLabV3 and SwinUNETR [13]. We also compared our method with two well-established models for prostate MRI segmentation (V-Net [30]) and liver CT segmentation (U-Net++ [54]), respectively.
@@ 3.4 Segmentation Accuracy
Table 1 reports the numerical results of all methods on the prostate MRI dataset. To investigate the statistical significance of the proposed method over compared methods on each metrics, the Wilcoxon signed-rank test was employed and also reported in Table 1. It can be observed that our CE-enhanced 2D networks can be regarded as significantly better than the conventional 2D networks on the DSC evaluation metric. Furthermore, regarding the backbones of U-Net and DeepLabV3, the $p$-values of CE-enhanced 2D networks $vs$. 3D networks on almost all metrics are beyond the 0.05 level, which indicates that the CE-enhanced 2D networks achieved similar performance as the 3D U-Net and 3D DeepLabV3 did. However, due to the superior performance of the cutting-edge 3D Transformer, the Missformer+CE could not attain similar segmentation results compared to the 3D SwinUNETR. Fig. 7 visualizes the segmentation results in transversal, coronal and sagittal views, respectively. Apparently, our CE-enhanced networks obtained similar segmented boundaries to the ground truth. Compared with conventional 2D networks, our method generated more accurate and continuous results in volumetric view (see the coronal and sagittal views in Fig. 7). In general, the results shown in Table 1 and Fig. 7 prove the effectiveness of our contextual embedding strategy. Compared with conventional 2D networks, the proposed CE block leveraged the learned embedding and the slice-wisely neighboring matching as an internal guidance to transfer contextual information thus facilitating 2D networks to capture volumetric representation properly.
Methods DSC (%) ASSD (mm) 95HD (mm) Sensitivity (%) Precision (%)
UNet++ [54] 94.23 $\pm$ 2.94* 0.95 $\pm$ 0.58* 4.89 $\pm$ 6.50* 96.05 $\pm$ 1.96 96.43 $\pm$ 1.39
U-Net [37] 94.13 $\pm$ 2.32* 1.84 $\pm$ 1.12* 8.62 $\pm$ 7.62* 96.14 $\pm$ 1.47 96.54 $\pm$ 1.08
3D U-Net [8] 95.67 $\pm$ 2.23* 2.31 $\pm$ 1.72* 2.31 $\pm$ 1.72 97.14 $\pm$ 0.94 96.38 $\pm$ 0.86
U-Net+CE (Ours) 96.49 $\pm$ 1.01 0.50 $\pm$ 0.17 1.77 $\pm$ 0.84 96.21 $\pm$ 1.52 96.75 $\pm$ 1.07
DeepLabV3 [6] 95.05 $\pm$ 4.85* 1.78 $\pm$ 2.24* 7.62 $\pm$ 14.83* 94.98 $\pm$ 1.30 95.84 $\pm$ 0.93
3D DeepLabV3 95.33 $\pm$ 1.44 1.49 $\pm$ 0.81* 5.83 $\pm$ 5.56* 94.91 $\pm$ 1.55 94.55 $\pm$ 1.29*
DeepLabV3+CE (Ours) 95.59 $\pm$ 2.03 0.61 $\pm$ 0.37 1.95 $\pm$ 1.50 95.88 $\pm$ 1.00 96.09 $\pm$ 0.74
Missformer [19] 96.45 $\pm$ 0.40 1.14 $\pm$ 0.20 3.81 $\pm$ 0.81 96.58 $\pm$ 1.39 96.36 $\pm$ 1.06
SwinUNETR [13] 96.63 $\pm$ 0.46 1.10 $\pm$ 0.20 3.24 $\pm$ 0.50 97.14 $\pm$ 1.43 96.17 $\pm$ 1.27
Missformer+CE (Ours) 96.56 $\pm$ 0.42 1.11 $\pm$ 0.15 3.55 $\pm$ 0.63 96.53 $\pm$ 1.47 96.62 $\pm$ 0.96
     Methods      ##\## Params      FLOPs (G)      Input size
     UNet [37]      7.76 M      3.42      (1, 1, 128, 128)
     3D UNet [8]      5.64 M      61.97      (1, 1, 64, 128, 128)
     U-Net+CE (Ours)      7.96 M      6.62      (1, 1, 128, 128)
     DeepLabV3 [6]      25.42 M      6.68      (1, 1, 128, 128)
     3D DeepLabV3      74.72 M      322.17      (1, 1, 64, 128, 128)
     DeepLabV3+CE (Ours)      25.63 M      7.66      (1, 1, 128, 128)
     Missformer [19]      10.77 M      6.95      (1, 1, 224, 224)
     3D Swin unetr [13]      15.7 M      12.71      (1, 1, 64, 64, 64)
     Missformer+CE (Ours)      10.97 M      27.36      (1, 1, 224, 224)
The quantitative results on the liver CT dataset are listed in Table 2. It can be observed that the CE-enhanced networks outperformed conventional 2D networks on all metrics. It is worth noting that the proposed CE block largely improved the 2D U-Net and DeepLabV3 with respect to the metrics of ASSD and 95HD, corroborating that the CE block has strong capability to utilize the slice-wise context for the accurate volumetric segmentation. The statistical results in Table 2 further shows our method can make the segmentation performance of the 2D networks close to or even surpass that of the 3D networks. Figs. 8 visualizes the liver segmentation results in transversal, coronal and sagittal views, respectively. Our method can successfully infer the ambiguous boundaries and attain an accurate and smooth volumetric segmentation.